OpenStudy (anonymous):

Rahki is making a cardboard cut out of a right triangle. The longest side must be 21 inches long, and one of the short sides must be 13 inches long. How many inches long will he make the other side? Use the appropriate formula to find the solution and round to the nearest inch. 20 in 16 in 14 in 25 in

OpenStudy (anonymous):

\[\sqrt{21^2-13^2}=4 \sqrt{17}, \text{ Round}\left[4 \sqrt{17},1\right]=16 \]

OpenStudy (anonymous):

Thank you for the medal.

OpenStudy (anonymous):

lol yw thanks for the help(: